OpenStudy (anonymous):

What is the solution of log4x − 616 = 4?

OpenStudy (anonymous):

@Navket_Jha my answer choices are -2, 2, 3, and 4.

OpenStudy (anonymous):

how did you solve it?

OpenStudy (anonymous):

@Navket_Jha

OpenStudy (anonymous):

log(4x) = 620, what properties of log can help?

OpenStudy (anonymous):

or is that log with base 4 of x?

OpenStudy (anonymous):

Check your question. I think something's missing

OpenStudy (anonymous):

Is it \[\log_{4} x\] or log4x?

OpenStudy (anonymous):

the first one you typed

OpenStudy (anonymous):

Ok so having log with a base of 4 you do the following: log4x-616=4 log4x=620 log4x=620 means x^620 = 4

OpenStudy (anonymous):

oeps

OpenStudy (anonymous):

I mean

OpenStudy (anonymous):

?

OpenStudy (anonymous):

4^620=x

OpenStudy (anonymous):

wtf.. wait a minute im confused

OpenStudy (anonymous):